OpenStudy (anonymous):

i dint get how to do this -3(t – 16) – 19 = 23

OpenStudy (cwrw238):

first expand the brackets (use distributive law) -3*t - (-3)*(-16) - 19 = 23 -3t + 48 - 19 = 23 now subtract 48 and add 19 to both sides of the '=' : -3t + 48 - 48 - 19 + 19 = 23 -48 + 19 -3t = -25 + 19 -3t = -6 divide both sides by -3 t = -6 /-3 t = 2

OpenStudy (anonymous):

thank you very much

OpenStudy (cwrw238):

* correction in first line -3*t -3 * (-16) - 19 = 23

OpenStudy (cwrw238):

the rest is ok

OpenStudy (cwrw238):

yw